# Quantumness certification via non-demolition measurements

### v1

@@ III Quantum non-demolition measurement
The function $\mathcal{G}_{\lambda}$ defines the QNDM quasi-characteristic function associated with the correlation function of $\hat{A}$ measured at $t_{0},t_{1}$. It can be shown, indeed, that the derivatives of $\mathcal{G}_{\lambda}$ with respect to $\lambda$ at $\lambda=0$ are proportional to the moments of the distribution of the $\hat{A}$ eigenvalues [20, 21, 31, 22, 23]. Writing $\mathcal{G}_{\lambda}$ as [19, 24, 25]
$\mathcal{G}_{\lambda}={\rm Tr}\left[e^{i\frac{\lambda}{2}\hat{A}}\hat{U}_{1}\,e^{i\frac{\lambda}{2}\hat{A}}\hat{\rho}^{0}e^{i\frac{\lambda}{2}\hat{A}}\hat{U}_{1}^{\dagger}\,e^{i\frac{\lambda}{2}\hat{A}}\right$]. (9)
and taking the first derivatives of $\mathcal{G}_{\lambda}$ with respect to $\lambda$ and setting $\lambda=0$, we obtain the expectation value of the variation of $\hat{A}$ within the time-interval $t_{0},t_{1}] [$19, 24]:
@@
where $\hat{\rho}(t_{1})=\hat{U}_{1}\hat{\rho}^{0}\hat{U}_{1}^{\dagger}$, $\langle\hat{A}(t_{1})\rangle={\rm Tr}[\hat{A}\hat{\rho}(t_{1})]$, and $\hat{A}^{H}(t_{1})\mathrel{\mathop{\ordinarycolon}}=\hat{U}_{1}^{\dagger}\hat{A}\hat{U}_{1}$ with the superscript $H$ denoting the Heisenberg representation. Moreover, by direct calculation [24], it can be shown that the second derivative of $\mathcal{G}_{\lambda}$ with respect to $\lambda$, computed at $\lambda=0$, is proportional to ${\rm Tr}\left[\left(\hat{A}^{H}(t_{1})-\hat{A}(t_{0})\right)^{2}\hat{\rho}^{0}\right]$, which is the second moment of the variation of $\hat{A}$.
@@
With these clarifications, given the final state in Eq. (7), we can calculate the QNDM quasi-characteristic function $\mathcal{G}_{\lambda}$
$\mathcal{G}_{\lambda}=\sum_{i,l,j}e^{i\lambda\left(a_{j}+\frac{a_{i}+a_{l}}{2}\right)}U_{1,ji}\rho^{0}_{il}U^{*}_{1,lj}\,$, (11)
that can be directly obtained experimentally as it is associated with the measurement of a phase, and, in this sense, it has a clear operational meaning.
Starting from the observation that $\mathcal{G}_{\lambda}$ is a quasi-characteristic function, it is natural to ask what kind of probability distribution can be associated with it. As usual, the corresponding probability distribution is obtained by taking the inverse Fourier transform of $\mathcal{G}_{\lambda}$, i.e., $\mathcal{P}_{\rm QNDM}(\Delta)=(2\pi)^{-1}\int\exp\{-i\lambda\Delta\}\mathcal{G}_{\lambda}d\lambda$. It turns out that, in general, $\mathcal{P}_{\rm QNDM}(\Delta)$ is not positively defined and, therefore, is a quasi-probability density function. We must stress that $\mathcal{P}_{\rm QNDM}(\Delta)$ cannot be directly measured in an experiment, but reconstructed from the manipulation of the experimental data. This means that the negative regions of $\mathcal{P}_{\rm QNDM}(\Delta)$ have a physical interpretation, as discussed below, but are not associated with any physical observable.
@@
It is interesting to notice that $P_{\rm QNDM}(j,i,l)$ contains the Kirkwood-Dirac one [25, 63], which can be obtained by summing over the index $l$ related to the quantum measurement with projectors $\{\hat{\Pi}_{l}\}$.
Figure 1: A pictorial representation of the system’s time evolution in terms of the outcomes from sequential measurements performed at three times $(t_{0},t_{1},t_{2})$. In this picture, the trajectories developed by the measured system are ascribable to Feynman paths built over the measurement outcomes. The possible outcomes of the measurements are $a_{i}$ and $a_{l}$ at time $t_{0}$, $a_{j}$ and $a_{m}$ at time $t_{1}$, and $a_{k}$ at time $t_{2}$. The sequence of outcomes identifies the paths that the monitored system could follow during its time evolution. The red dashed curve represents a classical (macrorealistic) path where, at any time, we have a single outcome (e.g., $a_{i}\rightarrow a_{m}\rightarrow a_{k}$) so that the observable has a determinate outcome. The blue curves present quantum paths in which the system is in a superposition of states with different outcomes. In this case, the statistics of outcomes got from measuring $\hat{A}$ change as the paths are no longer classical and violate the macrorealism requirement.
@@
Eq. (13) allows us to have a direct interpretation of the contributions. The function $P_{\rm QNDM}(j,i,l)$ is the probability amplitude to arrive in the state $\mathinner{|{j}\rangle}$ following the superposition of paths $i\rightarrow j$ and $l\rightarrow j$. A pictorial representation is presented in Fig. 1 where the measurement sequence is associated with a trajectory in the measurement outcomes space, e.g., $a_{i}\rightarrow a_{j}\rightarrow a_{k}$.
@@
Since $P_{\rm QNDM}(j,i,l)$ are probability amplitudes, they may take negative values from which the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ comes. The $P_{\rm QNDM}(j,i,l)$ can be even complex numbers with imaginary parts, but the latter do not contribute to the non-positivity of the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$ for symmetry reasons, as it will be clarified below.
@@
To understand the implications of the quasi-probability density function (13), let us separate its classical and quantum contributions. The classical contributions $\mathcal{P}_{\rm cl}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$ are the ones in which, during the whole evolution, the system is always in an eigenstate of the measured observable (see Fig. 1 for a pictorial representation); as a consequence, they satisfy the MRps condition as presented in Sec. I. These are the contributions that can be assessed through the TPM protocol that relies on sequential projective measurements. The latter correspond to trajectories that can be associated with a positive (multitime) probability, and thus ascribable to classical evolutions in statistical mechanics. Formally, they are obtained from $\mathcal{P}_{\rm QNDM}(\Delta)$ when $l=i$, i.e.,
$\mathcal{P}_{\rm cl}(\Delta)=\sum_{i,j}P_{\rm QNDM}(j,i,i)\delta\left[\Delta-\Delta_{j,i,i}\right],$ (15)
where
@@
On the other hand, the quantum contributions $\mathcal{P}_{\rm q}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$ are the ones for which the system is in a superposition of the eigenstates of $\hat{A}$ at any time $t$ of the whole evolution (see Fig. 1) and, in the terminology used in Sec. I and Ref. [5], they are the ones that violate the MRps assumption. In the easiest picture with two measurement times (e.g., $t_{0},t_{1}$), the quantum contributions originate from considering $l\neq i$, which reflects the fact that the system is initially in a coherent superposition of the eigenstates $\mathinner{|{i}\rangle}$ and $\mathinner{|{l}\rangle}$ of $\hat{A}$. Hence, formally,
@@
Let us now suppose that the measurement outcome $a_{i}$ can be only $\pm 1$. The classical contributions in Eq. (15) are associated with $\Delta_{j,i,i}=a_{j}+a_{i}$ that assume values 0 and $\pm 2$. On the contrary, the quantum contributions in Eq. (17) are associated with $\Delta_{j,i,l}$ with $l\neq i$ that also take values $\pm 1$. Therefore, in this sense, the quantum contributions of $\mathcal{P}_{\rm QNDM}(\Delta)$ are linked with sequences of measurement outcomes that are forbidden in classical transitions.
@@ III.1 Application to driven systems: Measurement of quantum work
The results in Sec. III can be applied to discuss how the QNDM approach can be exploited to measure the work in a closed quantum system. We only need to assume that at time $t=0$ the coupling between the system and the detector occurs with $\hat{A}=\hat{H}(0)$, at time $\mathcal{T}$ it is with $\hat{A}=\hat{H}(\mathcal{T})$ and, in the second coupling transformation, we flip the coupling sign $\lambda\rightarrow-\lambda$ to store the variation, i.e., the difference, of the system energy in the detector state during the composite evolution.
@@
The average work can be calculated from the derivative of the corresponding quasi-characteristic function with respect to $\lambda$, as shown above, or directly from the distribution (18). It turns out to be equal to the expected average in Eq. (4), which is not recovered by the TPM scheme whenever the initial density operator $\hat{\rho}^{0}$ is not diagonal along the eigenbasis of $\hat{H}(0)$ [47, 48].
Importantly, we have observed that every time the TPM protocol fails to reproduce the unperturbed value ${\rm Tr}[\hat{W}\hat{\rho}^{0}]$ for the average work $\mathcal{W}$, then the QNDM quasi-probability density function of the work exhibits at least a negative region. The reason for this behaviour has a broader explanation that we can identify in the violation of macrorealism, as we will discuss in the following. Specifically, for a work distribution, these negative regions are associated with non-classical energy exchanges. For example, consider a two-level system with bare Hamiltonian $\hat{H}=(\hbar\omega/2)\hat{\sigma}_{z}$, where $\hat{\sigma}_{z}$ is the Pauli matrix along the $z$-axis. The system is driven by a cyclic transformation. This means that the energy spectrum is the same at the beginning and at the end of the system time evolution, i.e., the energies $\epsilon_{j}(\mathcal{T})$ and $\epsilon_{j}(0)$ ($j=1,2$) take the same values $\pm\hbar\omega/2$. The TPM scheme — providing the classical contribution to $\mathcal{P}_{\rm QNDM}(W)$ — allows only for energy exchanges equal to 0 and $\pm\hbar\omega$. On the contrary, due to the presence of quantum coherence in the initial density operators $\hat{\rho}^{0}$, the work distribution develops quantum contributions, and the support of $\mathcal{P}_{\rm QNDM}(W)$ comprises the work realizations $\pm\hbar\omega/2$, 0, and $\pm\hbar\omega$. The ‘fractional’ energy exchanges $\pm\hbar\omega/2$ are not allowed by any classical and statistical interpretation, and cannot be recovered by the TPM scheme.
We point out that the experimental protocol consists in measuring experimentally the quasi-characteristic function $\mathcal{G}_{\lambda}$ such that $\mathcal{P}_{\rm QNDM}(W)$ is reconstructed from the analysis of the experimental data. In addition, we observe that the first and second moments of the QNDM work quasi-probability density function have a clear physical interpretation, extending the validity of these concepts from classical thermodynamics. However, the structure (in particular, the time-ordering) of the higher statistical moments may not be physically motivated. The latter, indeed, are ultimately determined and cannot be separated by the measurement procedure.
@@ III.2 Non-demolition measurements for open quantum system
In this section, we briefly discuss the application of the QNDM protocol to open quantum systems [68]. We will then make use of these concepts in Sec. VI, where they will be contextualized for the tracking of the quantum-to-classical transition.
In order to keep the presentation affordable, we are going to argue the main ideas and advantages of applying QNDM to monitor the dissipated heat in coherently-driven and open (due to system-environment interactions) dynamics.
The dissipated heat is a process function, and thus depends on the full time evolution of the system. For example, a quantum system starting and ending in the ground state does not change its overall internal energy. However, due to the interaction with the external environment, the system can produce a possibly arbitrary amount of heat during the open dynamics, through a sequence of excitations and relaxations. The only way to keep track of the dissipated heat is by constantly monitoring the system dynamics.
When it comes to the dissipated heat at the quantum level, a standard approach to this problem [69] is to perform sequential measurements of the environment degrees of freedom, in order to evaluate the energy dissipated into the environment, which is identified as heat. However, this approach may be both conceptually and practically unfeasible.
First, in many applications, a nanoscale system couples with an environment that has to be considered macroscopic, namely possessing a virtually infinite number of degrees of freedom and an essentially infinite heat capacitance. This entails a common approximation (Born’s approximation) [68, 70] done in dealing with open quantum system dynamics, i.e., to assume that the state of the environment does not change because of the interaction with the nanoscale system. As a result, from a practical point of view, one would need to detect any absorbed/emitted energy quantum from the environment, or at least to statistically sample a large portion of the environmental degrees of freedom.
@@
The QNDM approach can overcome these drawbacks for monitoring the heat, because it is designed to extract information from the system degrees of freedom, without inducing any perturbation on the dynamics (no-signalling in time condition). The important requirement to use the QNDM approach to monitor the heat over time is that the time-scales of the external drive (performing work on the system) and the environment are separated [20]. Under this assumption, we can exploit [19] the result stating that time-changes of the system Hamiltonian are associated with work, while time-changes in the system density operator are associated with dissipated heat [77]. Hence, performing fast sequential measurements, the monitoring of the system’s internal energy changes leads to reconstructing energy dissipation in terms of heat. Then, over a longer time scale, the effects due to the external drive in terms of work can also be included, with the constraint of respecting the first law of thermodynamics [20].
IV Macrorealism, quasi-probability distribution and Leggett-Garg inequalities
IV.1 Macrorealism
The original definition of Macrorealism per se (MRps) [5], which was prone to different interpretations, has been recently clarified [78, 79]. Specifically, once the observable to measure is identified, the assumption of MRps is satisfied if quantum superpositions are not allowed, at any time, within the Hilbert space spanned by the eigenbasis of the observable or by a statistical mixture of them [79]. The assumption of MRps is thus satisfied if there cannot be coherent superpositions and the state of the system is always diagonal in the basis that diagonalizes the measurement observable [79, 80].
The Noninvasive Measurability (NIM) assumption, entering the Macrorealism (MR), is equally difficult to formalize. For this reason, it is usually flanked by the No-Signalling In Time (NSIT) conditions [7, 6] that are a less strong version of NIM. This allows us to relate the NIM with the possibility of attaining the marginal probability of single events from the joint probability of a sequence of them, without the results being perturbed by any intermediate measurement [6]. As an example, if $p(s(t_{0}),s(t_{1}),s(t_{2}))$ is a generic joint probability at the three times $(t_{0},t_{1},t_{2})$, then the NSIT condition for the outcome $s(t_{2})$ ensures that $\sum_{s(t_{0}),s(t_{1})}p(s(t_{0}),s(t_{1}),s(t_{2}))=P(s(t_{2}))$, where $P(s(t_{2}))$ is the probability to measure $s(t_{2})$ in correspondence of the system state at time $t_{2}$ where the measurement is performed. This assures that no previous measurements in the sequence have statistically affected the system state. Throughout the remaining part of the paper, the symbol $P$ will denote both probabilities at single times, given by Born’s rule, and probabilities at multiple times respecting the theory of classical probability (in particular, the Kolmogorov axioms).
Below, we will show that the QNDM quasiprobabilities always satisfy the NSIT assumption; hence, the violation of MR can be reduced only to the violation of the MRps, for which both a necessary and sufficient condition can be given. Formally, ensuring the validity of such a necessary and sufficient condition certifies the presence and/or development of quantum features in the system state under scrutiny.
IV.2 Three-measurement scheme

### v2

@@ III Quantum non-demolition measurement
The function $\mathcal{G}_{\lambda}$ defines the QNDM quasi-characteristic function associated with the correlation function of $\hat{A}$ measured at times $t_{0},t_{1}$. It can be shown, indeed, that the derivatives of $\mathcal{G}_{\lambda}$ with respect to $\lambda$ at $\lambda=0$ are proportional to the moments of the distribution built with the eigenvalues of $\hat{A}$ [20, 21, 42, 22, 23]. Writing $\mathcal{G}_{\lambda}$ as [19, 24, 25]
$\mathcal{G}_{\lambda}={\rm Tr}\left[e^{i\frac{\lambda}{2}\hat{A}}\hat{U}_{1}\,e^{i\frac{\lambda}{2}\hat{A}}\hat{\rho}^{0}e^{i\frac{\lambda}{2}\hat{A}}\hat{U}_{1}^{\dagger}\,e^{i\frac{\lambda}{2}\hat{A}}\right$], (9)
taking the first derivatives of $\mathcal{G}_{\lambda}$ with respect to $\lambda$ and setting $\lambda=0$, we obtain the expectation value of the variation of $\hat{A}$ within the time-interval $t_{0},t_{1}] [$19, 24]:
@@
where $\hat{\rho}(t_{1})=\hat{U}_{1}\hat{\rho}^{0}\hat{U}_{1}^{\dagger}$, $\langle\hat{A}(t_{1})\rangle={\rm Tr}[\hat{A}\hat{\rho}(t_{1})]$ and $\hat{A}^{H}(t_{1})\mathrel{\mathop{\ordinarycolon}}=\hat{U}_{1}^{\dagger}\hat{A}\hat{U}_{1}$, with the superscript $H$ denoting the Heisenberg representation. Moreover, by direct calculation [24], it can be shown that the second derivative of $\mathcal{G}_{\lambda}$ with respect to $\lambda$, computed at $\lambda=0$, is proportional to ${\rm Tr}\left[\left(\hat{A}^{H}(t_{1})-\hat{A}(t_{0})\right)^{2}\hat{\rho}^{0}\right]$, which is the second moment of the variation of $\hat{A}$ at times $t_{0},t_{1}$.
@@
With these clarifications and given the final state in Eq. (7), the expression of the QNDM quasi-characteristic function $\mathcal{G}_{\lambda}$ simplifies to
$\mathcal{G}_{\lambda}=\sum_{i,l,j}e^{i\lambda\left(a_{j}+\frac{a_{i}+a_{l}}{2}\right)}U_{1,ji}\rho^{0}_{il}U^{*}_{1,lj}$ (11)
The right-hand side of Eq. (11) can be directly obtained experimentally, as it is associated with protocols measuring a phase; in this sense, it has a clear operational meaning.
Starting from the observation that $\mathcal{G}_{\lambda}$ is a quasi-characteristic function, it is natural to ask what kind of probability distribution can be associated with it. As usual, the corresponding probability distribution is obtained by taking the inverse Fourier transform of $\mathcal{G}_{\lambda}$, i.e., $\mathcal{P}_{\rm QNDM}(\Delta)=(2\pi)^{-1}\int\exp\{-i\lambda\Delta\}\mathcal{G}_{\lambda}d\lambda$. It turns out that, in general, $\mathcal{P}_{\rm QNDM}(\Delta)$ is not positively defined and, therefore, is a quasi-probability density function. We must stress that $\mathcal{P}_{\rm QNDM}(\Delta)$ can be reconstructed experimentally; this means that the negative regions of $\mathcal{P}_{\rm QNDM}(\Delta)$ have a physical interpretation, as discussed below, albeit they are not associated with any physical observable.
@@
It is important to note that $P_{\rm QNDM}(j,i,l)$ contains the Kirkwood-Dirac one [25, 68], which is obtained by summing over the index $l$ related to the quantum measurement with projectors $\{\hat{\Pi}_{l}\}$.
@@
In Eq. (13), the function $P_{\rm QNDM}(j,i,l)$ is the interference contribution between two alternative and superposed paths $i\rightarrow j$ and $l\rightarrow j$. A pictorial representation is presented in Fig. 1 where the measurement sequence is associated with a trajectory in the measurement outcomes space, e.g., $a_{i}\rightarrow a_{j}\rightarrow a_{k}$.
Figure 1: Pictorial representation of the system’s time evolution in terms of the outcomes from the sequential measurement at the three times $(t_{0},t_{1},t_{2})$. In this picture, the trajectories developed by the measured system are ascribable to paths built over the measurement outcomes. The possible outcomes of the measurements are $a_{i}$ and $a_{l}$ at time $t_{0}$, $a_{j}$ and $a_{m}$ at time $t_{1}$, and $a_{k}$ at time $t_{2}$. The sequence of outcomes identifies the paths that the monitored system could follow during its time evolution. The red dashed curve represents a classical path where, at any time, we have a single outcome (e.g., $a_{i}\rightarrow a_{m}\rightarrow a_{k}$) so that the observable has a determinate outcome. The blue curves present quantum paths in which the system is in a superposition of states with different outcomes. In such a case, the statistics of outcomes obtained from measuring $\hat{A}$ at times $(t_{0},t_{1},t_{2})$ change as there is a coherent superposition of paths that violates the requirement of macrorealism per se.
@@
Since $P_{\rm QNDM}(j,i,l)$ are interference contributions, they may take negative values from which the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ comes. The $P_{\rm QNDM}(j,i,l)$ can be even complex numbers with imaginary parts, but the latter do not contribute to the non-positivity of the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$ for symmetry reasons [25].
@@
To understand the implications of the quasi-probability density function (13), let us separate its classical and quantum contributions. The classical contributions $\mathcal{P}_{\rm cl}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$ are the ones in which, during the whole evolution, the system is always in an eigenstate of the measured observable (see Fig. 1 for a pictorial representation); as a consequence, they satisfy the MRps condition presented in the Introduction (Sec. I). These are the contributions that can be assessed through the TPM protocol that relies on sequential projective measurements. The latter correspond to trajectories associated with a positive (multitime) probability, and thus ascribable to classical evolutions in statistical mechanics. Formally, they are obtained from $\mathcal{P}_{\rm QNDM}(\Delta)$ when $l=i$, i.e.,
$\mathcal{P}_{\rm cl}(\Delta)=\sum_{i,j}P_{\rm QNDM}(j,i,i)\delta\left[\Delta-\Delta_{j,i,i}\right]$ (15)
with
@@
On the other hand, the quantum contributions $\mathcal{P}_{\rm q}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$ are the ones for which the system is in a superposition of the eigenstates of $\hat{A}$ at any time $t$ of the whole evolution (see blue curves in Fig. 1) and, in the terminology used in Sec. I and Ref. [5], they are the ones that violate the MRps assumption. In the easiest picture with two measurement times (e.g., $t_{0},t_{1}$), the quantum contributions originate from considering $l\neq i$, which reflects the fact that the system is initially in a coherent superposition of the eigenstates $\mathinner{|{i}\rangle}$ and $\mathinner{|{l}\rangle}$ of $\hat{A}$. Hence, formally,
@@
Let us now suppose that the measurement outcome $a_{i}$ can be only $\pm 1$. The classical contributions in Eq. (15) are associated with $\Delta_{j,i,i}=a_{j}+a_{i}$ that assume the values 0 and $\pm 2$. On the contrary, the quantum contributions in Eq. (17) are associated with $\Delta_{j,i,l}$ with $l\neq i$ that also take the values $\pm 1$. Therefore, in this sense, the quantum contributions of $\mathcal{P}_{\rm QNDM}(\Delta)$ are linked with sequences of measurement outcomes that are forbidden in classical transitions.
@@ III.1 Application to driven systems: Measurement of quantum work
The results in Sec. III can be applied to discuss how the QNDM approach can be exploited to measure the work in a closed quantum system. To do this, we need to assume that at time $t=0$ the coupling between the system and the detector occurs with $\hat{A}=\hat{H}(0)$, and at time $\mathcal{T}$ it is with $\hat{A}=\hat{H}(\mathcal{T})$. Moreover, in the second coupling transformation, we have to flip the coupling sign $\lambda\rightarrow-\lambda$, to store the variation of the system energy in the detector state during the composite evolution.
@@
The average work can be calculated from the derivative of the corresponding quasi-characteristic function with respect to $\lambda$, or directly from the distribution (18). It turns out it is equal to the expected average in Eq. (4), which is not recovered by the TPM scheme whenever the initial density operator $\hat{\rho}^{0}$ is not diagonal along the eigenbasis of $\hat{H}(0)$ [53, 54]. Importantly, we have observed that every time the TPM protocol fails to reproduce the unperturbed value ${\rm Tr}[\hat{W}\hat{\rho}^{0}]$ for the average work $\mathcal{W}$, then the QNDM quasi-probability density function of the work exhibits at least a negative region. Specifically, for a work distribution, these negative regions are associated with non-classical energy exchanges. For example, consider a two-level system with bare Hamiltonian $\hat{H}=(\hbar\omega/2)\hat{\sigma}_{z}$, where $\hat{\sigma}_{z}$ is the Pauli matrix along the $z$-axis. The system is driven by a cyclic transformation. This means that the energy spectrum is the same at the beginning and at the end of the system time evolution, i.e., the energies $\epsilon_{j}(\mathcal{T})$ and $\epsilon_{j}(0)$ ($j=1,2$) take the same values $\pm\hbar\omega/2$. The TPM scheme — providing the classical contribution to $\mathcal{P}_{\rm QNDM}(W)$ — allows only for energy exchanges equal to 0 and $\pm\hbar\omega$. On the contrary, due to the presence of quantum coherence in the initial density operators $\hat{\rho}^{0}$, the work distribution develops quantum contributions, and the support of $\mathcal{P}_{\rm QNDM}(W)$ comprises the work realizations $\pm\hbar\omega/2$, 0, and $\pm\hbar\omega$. The ‘fractional’ energy exchanges $\pm\hbar\omega/2$ are not allowed by any classical and statistical interpretation, and cannot be recovered by the TPM scheme.
We point out that the experimental protocol consists of measuring experimentally the quasi-characteristic function $\mathcal{G}_{\lambda}$ such that $\mathcal{P}_{\rm QNDM}(W)$ is reconstructed from the analysis of the experimental data. In addition, we observe that the first and second moments of the QNDM work quasi-probability density function have a clear physical interpretation, extending the validity of these concepts from classical thermodynamics. However, the structure (in particular, the time-ordering) of the higher statistical moments may not be physically motivated. The latter, indeed, are ultimately determined and cannot be separated by the measurement procedure.
@@ III.2 Non-demolition measurements for open quantum system
In this section, we briefly discuss the application of the QNDM protocol to open quantum systems [73]. We will then make use of these concepts in Sec. VI, where they are contextualized for the tracking of the quantum-to-classical transition. In order to keep the presentation affordable, here we just argue the main ideas and advantages of applying QNDM to monitor the dissipated heat in coherently-driven and open dynamics due to system-environment interactions.
The dissipated heat is a process function, and thus depends on the full time evolution of the system. For example, a quantum system starting and ending in the ground state does not change its overall internal energy. However, due to the interaction with the external environment, the system can produce a possibly arbitrary amount of heat during the open dynamics, through a sequence of excitations and relaxations. The only way to keep track of the dissipated heat is by constantly monitoring the dynamics of the system.
When it comes to the dissipated heat at the quantum level, a standard approach [74] is to perform sequential measurements of the environmental degrees of freedom, in order to evaluate the energy dissipated into the environment, which is identified as heat. However, this approach may be both conceptually and practically unfeasible.
First, in many applications, a nanoscale system couples with an environment that has to be considered macroscopic, namely possessing a virtually infinite number of degrees of freedom and an essentially infinite heat capacitance. This entails a common approximation (Born’s approximation) [73, 75] in open quantum systems, i.e., to assume that the state of the environment does not change because of the interaction with the nanoscale system. As a result, from a practical point of view, one would need to detect any absorbed/emitted energy quantum from the environment, or at least to statistically sample a large portion of the environmental degrees of freedom.
@@
The QNDM approach can overcome these drawbacks for monitoring the heat, because it is designed to extract information from the system’s degrees of freedom, without inducing any perturbation on the dynamics. The important requirement to use the QNDM approach for monitoring the heat over time is that the time-scales of the external drive (performing work on the system) and the environment are separated [20]. Under this assumption, we can exploit [19] the result stating that time-changes of the system’s Hamiltonian are associated with work, while time-changes in the density operator of the system are associated with dissipated heat [82]. Hence, performing fast system-detector couplings, the monitoring of the system’s internal energy changes leads to reconstructing energy dissipation in terms of heat. Then, over a longer time scale, the effects due to the external drive in terms of work can also be included, with the constraint of respecting the first law of thermodynamics [20].
IV Certification of quantum and classical behaviours
We are going to show that the negativity of the QNDM quasi-probability density function is a necessary and sufficient condition for the identification of quantum superpositions in the system’s state at the measurement times. Accordingly, given an initial system’s state and a measurement observable, determining a negative QNDM quasi-probability certifies whether the system under scrutiny exhibits quantum behaviours.
Furthermore, we will show that if the QNDM quasi-probability density function turns out to be positive, then the system has classical properties as intended in Ref. [25]; that is, the system’s states at the measurement times $t_{0},t_{1},t_{2}$ are necessarily diagonal with respect to the basis that spectrally decomposes the measurement observable. The latter condition entails that the QNDM quasi-probabilities can be provided by the TPM scheme (sequential projective measurements, operationally), meaning that it is identically equal to a classical probability distribution.
In order to effectively illustrate these findings, we discuss the prototypical case of three non-demolition measurements that allows us to compare the certification of quantum behaviours obtained from the QNDM and the standard LGIs.
IV.1 Three-measurement scheme